OpenStudy (rational):

solve \(\sin^5x + \cos^5x=1\) \(x\) is in first quadrant

OpenStudy (rational):

Easy to see that \(x=0,\pi/2\) are trivial solutions and wolfram gives http://www.wolframalpha.com/input/?i=solve+sin%5E5x%2Bcos%5E5x%3D1%2C+0%3Cx%3Cpi%2F2 idk how to prove that solutions don't exist for \(0\lt x\lt \pi/2\) ..

OpenStudy (anonymous):

looks cool http://www.wolframalpha.com/input/?i=x^2%2By^2%3D1%2Cx^5%2By^5%3D1

OpenStudy (rational):

that looks cool indeed xD i know it boils down to \(\sin^2x+\cos^2x=1\) somehow

OpenStudy (anonymous):

that is your only real condition right?

OpenStudy (rational):

if you mean x is real, yes..

OpenStudy (anonymous):

you are solving \[x^5+y^5=1\] given that \[x^2+y^2=1\]

OpenStudy (rational):

Ahh right! both problems are equivalent

OpenStudy (rational):

and i feel equally hard/easy

OpenStudy (anonymous):

yeah and i wouldn't bet more than $7 that there is not some more snappy approach

OpenStudy (anonymous):

maybe just graph \(x^5+y^5=1\)

OpenStudy (rational):

I think so... is it sufficient if we could show that \(x^5+y^5\lt x^2+y^2\) for \(x,y\in (0,1)\)

OpenStudy (rational):

Wow! that does it! thnks @satellite73

OpenStudy (rational):

sure calculus can solve everything :) it is easy to notice that \(a^n \lt a^m\) for \(a\in (0, 1)\) and \(n\gt m\)

OpenStudy (rational):

its not stupid, i think its an alternative approach...

OpenStudy (rational):

Here is the complete proof : for \(x\in(0,1)\) we have \(\sin^5x\lt \sin^2x\) and \(\cos^5x\lt \cos^2x\). that implies \(\sin^5x+\cos^5x \lt \sin^2x+\cos^2x=1\) \(\blacksquare\)

OpenStudy (ikram002p):

that solves it